OpenStudy (coconut):

what is the simplified form of (y^2 + 7y + 12) / (y^2 -2y - 15

OpenStudy (anonymous):

we can factor the top to (y+4)(y+3) and the denominator by (y-5)(y+3) so now we have (y+4)(y+3)/(y-5)(y+3) = (y+4)/(y-5)

OpenStudy (anonymous):

(y^2 + 7y + 12) / (y^2 -2y - 15 =(y+3)(y+4)/(y-5)(y+3) cancel (y+3) =(y+4)/(y-5) ans...